OpenStudy (anonymous):

find two consecutive whole numbers that sqrt 52lies between

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... all it's asking is the integers BEFORE and AFTER the decimal value of \(\bf \sqrt{52}\) so the numbers "a" and "b" \(\large {\color{purple}{ a}}...\sqrt{52}...{\color{purple}{ b}}\)

OpenStudy (anonymous):

like what times what makes 52?

OpenStudy (jdoe0001):

nope, just the "whole numbers" or integers that lie before and after \(\bf \sqrt{52}\)

OpenStudy (anonymous):

like 6 and 8?

OpenStudy (jdoe0001):

 well, consecutive, thus